# RACS and SADL: Towards Robust SMR in the Wide-Area Network

### v2

@@ Abstract
This paper proposes RACS, a novel randomized consensus protocol that ensures robustness against adversarial network conditions. RACS achieves optimal one-round trip latency under synchronous network conditions while remaining resilient to adversarial network conditions. RACS follows a simple design inspired by Raft, the most widely used consensus protocol in the cloud, and therefore enables seamless integration with the existing cloud software stack— a goal no previous asynchronous protocol has successfully achieved.
Experiments with a prototype deployed on Amazon EC2 confirm that RACS achieves a throughput of 28k cmd/sec under adversarial cloud network conditions, whereas existing leader-based protocols such as Multi-Paxos and Raft provide less than 2.8k cmd/sec. Under synchronous network conditions, RACS matches the performance of Multi-Paxos and Raft, achieving a throughput of 200k cmd/sec with a latency of 300ms, confirming that RACS introduces no unnecessary overhead. Finally, SADL-RACS —an optimized version of RACS designed for high performance and robustness—achieves an impressive throughput of 500k cmd/sec under synchronous network conditions and 196k cmd/sec under adversarial network conditions, further enhancing both performance and robustness.
@@
consensus, adversarial, networks.
@@ I Introduction
In practice, leader-based protocols and their multi-leader extensions [20][21][22][23] are designed to detect adversarial network conditions using timeouts [14]. If the leader replica does not respond within a pre-configured timeout, the remaining replicas elect a new leader replica. Electing a new leader replica often involves a separate ”view change” sub routine [14], where each replica votes on a potential next leader replica.
@@
Asynchronous consensus protocols [25][26][27] ensure high robustness and performance under adversarial network conditions by leveraging randomization instead of relying on view change. However, to the best of our knowledge, asynchronous consensus protocols have never been deployed in cloud applications due to three main drawbacks: (1) they require significantly higher bandwidth compared to leader-based protocols [26]; (2) their design is often incompatible with the existing cloud software stack, hindering seamless integration and rapid adoption [27, 28]; and (3) their complexity makes them difficult to understand and implement, which limits industry adoption. Hence, most cloud providers continue to deploy leader-based consensus protocols despite their limited robustness against adversarial network conditions [2].
Problem Statement: Can we design a consensus protocol that (1) remains robust under adversarial network conditions, (2) achieves performance comparable to leader-based protocols in synchronous network conditions, and (3) maintains a simple design that allows easy integration with the existing cloud software stack?
@@
Achieving robustness against adversarial network conditions has already been explored [26, 28, 25, 27], and we explicitly state that this is ”not” our primary contribution. Instead, RACS, our novel protocol, simultaneously achieves the three goals G1, G2, and G3—an accomplishment that no existing asynchronous protocol has achieved before. Designing RACS while achieving G1, G2 and G3 simultaneously is not trivial. First, designing a protocol that ensures robustness against adversarial network conditions without incurring additional overhead under synchronous conditions is a non-trivial algorithmic challenge. Second, designing a robust protocol while ensuring ease of understanding and compatibility with existing cloud software stacks is a challenge that no previous asynchronous consensus protocol has successfully addressed.
@@
This paper makes the following key contributions:
@@
  • We provide the formal proofs of RACS.
@@ II-A Threat Model and Assumptions
Due to the FLP impossibility result [33], any deterministic algorithm cannot solve consensus under asynchrony even under a single replica failure. In RACS, we circumvent the FLP impossibility result using randomization. In a typical cloud network, there are periods in which the network behaves synchronously, followed by phases where the network shows adversarial behavior. RACS makes use of this network behavior and operates in two modes: (1) synchronous mode and (2) fallback mode. During the synchronous periods, RACS employs a leader-based design to reach consensus using one round trip network delay, and during the fallback mode, RACS employs randomization.
@@ II-B Consensus
Multi-leader protocols: Multi-leader protocols [21, 20, 23, 24] mitigate the leader bottleneck issue present in single-leader-based protocols. However, they still depend on view change subroutines to manage replica failures and network connectivity issues. Hence, under adversarial network conditions, multi-leader protocols also suffer from loss of performance. Other extensions of leader-based protocols [35][36][22][37] also lose liveness under adversarial network conditions due to their reliance on view change mechanisms. In contrast, our proposal RACS maintains liveness under adversarial network conditions.
Randomized consensus protocols: Randomized consensus protocols [38][25][27][26][39] overcome the adversarial network challenge by using randomization instead of view changes. However, they are rarely used in cloud applications due to (1) high performance overhead in synchronous network conditions, (2) incompatibility with the existing cloud software stack, and (3) complexity in understanding and implementing. In contrast, our protocol RACS provides a practical solution to adversarial network challenges by achieving: (1) robustness in adversarial network conditions, (2) optimal one round-trip commit latency in synchronous networks, and (3) a simple design for fast implementation and seamless industry adoption.
@@ III-C RACS Algorithm
Synchronous mode: The synchronous mode of RACS is a leader-based and is similar to Raft, as depicted in lines 1–11 of Figure 4.. The synchronous mode leader $L_{vcur}$ for each view $v$ is predetermined and known to all replicas on bootstrap, for example $L_{vcur}=v_{cur}\%n$.
@@
Fallback mode: The number of blocks proposed during the fallback mode is a tunable parameter that we explain in Section IV-A. For simplicity of illustration, we use 2 fallback blocks per view in the following description. Upon receiving $n/2+1$ $<$timeout$>$ messages, RACS enters the fallback mode of operation. In the fallback mode, all replicas act as leaders, concurrently. Each replica takes the highest $block_{high}$ they are aware of, forms a level $1$ fallback block $B$ with a monotonically increasing rank compared to the highest $block_{high}$ it received and sends a $<$propose-async$>$ message (see line 12–16 of Figure 4).
@@
Once $n/2+1$ replicas have submitted $<$asynchronous-complete$>$ messages, the fallback mode is complete. The remaining task is to consistently select one chain from the set of fallback chains. This step is crucial for maintaining agreement. If two replicas select different fallback chains, the system loses the guarantee that all replicas commit the same sequence of blocks. To ensure that each replica consistently chooses the same chain, we employ a common-coin, an existing technique [26, 39, 25] used in distributed algorithms.
@@
Upon receiving $n/2+1$ $<$asynchronous-complete$>$ messages, each replica flips a common-coin (see line 21 of Figure 4) to get the elected leader Lelected. Each replica commits a level $2$ fallback block $B$ from Lelected if $B$ arrived amongst the first $n/2+1$ $<$asynchronous-complete$>$ messages (see line 22 of Figure 4). If a replica observes that the level $2$ fallback block from Lelected does not appear in the first $n/2+1$ $<$asynchronous-complete$>$ messages, but appears in Bfall[Lelected], then, the replica sets blockhigh to Bfall[Lelected] (see line 23 of Figure 4). If a replica observes that the level 2 fallback block from Belected is not in the first $n/2+1$ $<$asynchronous-complete$>$ messages and Bfall[Lelected] is null, then the replica doesn’t update its blockhigh and blockcommit. All replicas then exit the fallback mode and resume the synchronous path by uni-casting a $<$vote$>$ message to the synchronous leader Lvcur of the next view (see line 27 of Figure 4).
RACS protocol summary: Based on the above explanation of the RACS protocol, its outcomes can be summarized as follows. First, by committing new commands during the fallback mode, RACS ensures robustness against adversarial network conditions. Second, RACS commits new client commands in one network round trip, and hence provides optimal performance under synchronous normal case conditions. Finally, since RACS follows a simple design and extends Raft, it enables seamless integration with the existing cloud software stack.
@@ III-D Correctness and Complexity
We now give the proof intuitions of RACS. The formal proofs have been deferred to the anonymous online appendix 111online appendix link due to page limit.
RACS only commits blocks proposed by replicas, thereby inherently satisfy the validity property. Because RACS does not commit the same block more than once, it trivially satisfies the integrity property. In the following sections, we focus on the agreement and termination properties.
@@
Proof: A block $B$ can be committed in 2 instances; (1) synchronous mode and (2) fallback mode.
In the synchronous mode, a block is committed by Lvcur when a majority of replicas vote for the synchronous block in the same round. Hence, if the leader commits a block $B$, then it is guaranteed that at least a $n/2+1$ majority of the replicas have $B$ as their $block_{high}$, or have $B$ in the causal history of $block_{high}$. Due to quorum intersection, all the future blocks will extend $B$ because each new block extends blockhigh.
In the fallback mode, the only fallback block $B$ of level $2$ sent by Lelected is committed if it is among the first $n/2+1$ $<$asynchronous-complete$>$ messages received in that view, which implies that a majority of the replicas have received and saved $B$ in Bfall[Lelected]. Hence if at least one node commits $B$ in the fallback mode, then it is guaranteed that at least a $n/2+1$ number of the replicas set $B$ as $block_{high}$, thus extending $B$ in the next round.
Therefore, we claim that agreement property holds in RACS.
Termination property: RACS ensures termination under both synchronous and adversarial network conditions.
Proof: During the synchronous mode, commands are committed in each round by the leader after receiving $<$votes$>$ from $n/2+1$ replicas, thereby satisfying the termination property.
In the fallback mode, if more than $n/2+1$ replicas enter the fallback mode, i.e. there are less than $n/2$ replicas on the synchronous mode, i.e. the synchronous path cannot progress anymore, therefore all replicas eventually enter the fallback mode. The fallback mode (with at least $n/2+1$ correct replicas in it) will eventually reach a point where $n/2+1$ correct replicas have sent $<$asynchronous-complete$>$ messages. The common-coin($v$) will therefore take place. If the result of the common-coin($v$) lands on one of the first $n/2+1$ replicas to have submitted an $<$asynchronous-complete$>$ message, then a new block is committed. Since the coin is unbiased, there is a probability $p$$>$$\frac{1}{2}$ that a new fallback block is committed in a given fallback mode of execution. Each replica therefore commits a new block in under 2 views in expectation and commits eventually with probability 1, thus ensuring the protocol’s termination.
@@ IV-A Tuning fallback blocks
In Section III, we presented the RACS protocol using exactly two blocks in the fallback mode. In practice, adversarial network conditions can persist for unpredictable periods, making the use of a fixed number of blocks per view in fallback mode suboptimal.
In RACS, we dynamically adjust the number of blocks in fallback mode using Multi-Armed Bandit techniques [43] to optimize performance under varying network conditions. We initially run RACS with 2 blocks per fallback mode, and each replica monitors s: the number of times the algorithm switches between synchronous mode and the fallback mode. A high s implies a network where adversarial conditions persist, whereas a low s indicates a transient network problem that has self-healed. RACS employs a Multi-Armed Bandit-based explore/exploit strategy, increasing the number of blocks per fallback mode when s is high and decreasing it when s is low. RACS uses the same technique as Raft’s reconfiguration to consistently configure s across all replicas. Further details of this approach are beyond the scope of this paper and are omitted due to space constraints.
@@ IV-B Application specific optimizations
IV-B1 Optimization for class 1 applications
To reduce latency for class 1 applications, we employ pipelining, a well-established technique in distributed algorithms [25, 13, 20]. In pipelined-RACS, the synchronous mode leader $L_{\text{vcur}}$ proposes multiple blocks before receiving $n/2+1$ $<$vote$>$ messages. This allows multiple blocks to be committed within a single round-trip network delay during synchronous network periods, thereby reducing commit latency.
IV-B2 Optimization for class 2 applications
@@
RACS blocks contain two types of data: (1) metadata about the block, including $block_{high}$, $block_{commit}$, $v$, and $r$, and (2) a batch of client commands. Typically, the metadata is only a few bytes in size, while the batch of client commands can span several kilobytes. Bundling both metadata and client commands within RACS blocks increases the block size, which limits the system’s throughput in synchronous mode due to the bandwidth bottleneck at the leader replica. This is a significant limitation of RACS, particularly for class 2 applications that demand high throughput.
To address this limitation, we propose SADL, a novel command dissemination layer that decouples command dissemination from the critical path of the consensus protocol. SADL is a consensus-agnostic command dissemination layer designed to reliably distribute client commands to at least a majority of replicas, ensuring efficient and fault-tolerant command propagation. Once SADL replicates commands across a majority of replicas, RACS layer no longer needs to include the commands in it’s blocks. Instead, RACS only embeds a small, fixed-size ”reference” to the batch already disseminated by SADL. Using a fixed-size reference instead of the entire command batch significantly reduces the RACS leader replica’s bandwidth consumption, leading to a substantial improvement in throughput, as demonstrated in Section VI.
Decoupling command dissemination from consensus has already been explored in the context of blockchain protocols [28, 47]. However, our SADL design significantly differs from existing approaches. Unlike existing decoupling mechanisms [28, 47], where the dissemination layer advances in a lock-step fashion, SADL allows each replica to build its own chain of SADL blocks independently of the pace of other replicas. This key difference enables SADL to achieve higher throughput than existing approaches. We present this subtle yet non-trivial distinction as a novel systems contribution of this paper.
@@ V Implementation
We implemented pipelined-RACS and SADL-RACS using Go version 1.18 [29], in 3661 and 4631 lines of codes, respectively, as counted by CLOC [48]. We use the Go network library and TCP [30] for reliable FIFO point-to-point links between replicas. We used Protobuf encoding [49] for message serialization.
Both RACS and SADL-RACS implement batching in both clients and replicas as in existing implementations of Multi-Paxos, Raft[42], and EPaxos [50].

### v3

# RACS-SADL: Robust and Understandable Randomized Consensus in the Cloud

@@ Abstract
We propose RACS, a novel randomized consensus protocol that ensures robustness against adversarial network conditions. RACS achieves optimal one-round trip latency under synchronous network conditions while remaining resilient to adversarial network conditions. RACS follows a simple design inspired by Raft, the most widely used consensus protocol in the cloud, and therefore enables seamless integration with the existing cloud software stack.
Experiments with a prototype running on Amazon EC2 show that RACS achieves 28k cmd/sec throughput, ninefold higher than Raft under adversarial cloud network conditions. Under synchronous network conditions, RACS matches the performance of Multi-Paxos and Raft, achieving a throughput of 200k cmd/sec with a median latency of 300ms, confirming that RACS introduces no unnecessary overhead. Finally, SADL-RACS, a throughput-optimized version of RACS, achieves a throughput of 500k cmd/sec, delivering 150% higher throughput than Raft.
@@
consensus, adversarial, networks, scalability.
@@ I Introduction
In practice, leader-based protocols and their multi-leader extensions [20][21][22][23] are designed to detect adversarial network conditions using timeouts [14]. If the leader replica does not respond within a pre-configured timeout, the remaining replicas elect a new leader replica. Electing a new leader replica often involves a separate ”view change” sub routine [14], where each replica votes for a potential next leader replica.
@@
Asynchronous consensus protocols [26][27] ensure high robustness and performance under adversarial network conditions by leveraging randomization instead of relying on view change. However, to the best of our knowledge, asynchronous consensus protocols have never been deployed in cloud applications due to three main drawbacks: (1) they require significantly higher bandwidth compared to leader-based protocols [26]; (2) their design is often incompatible with the existing cloud software stack, hindering seamless integration and rapid adoption; and (3) their complexity makes them difficult to understand and implement, which limits industry adoption. Hence, cloud providers continue to deploy leader-based consensus protocols despite their limited robustness against adversarial network conditions [2].
Problem Statement: Can we design a consensus protocol that (1) remains robust under adversarial network conditions, (2) achieves performance comparable to leader-based protocols under synchronous network conditions, and (3) maintains a simple design that allows easy integration with the existing cloud software stack?
@@
Achieving robustness against adversarial network conditions has already been explored [26, 28, 25, 27], and we explicitly state that this is ”not” our primary contribution. Instead, RACS, our novel protocol, ”simultaneously” achieves the three goals G1, G2, and G3—an accomplishment that no existing asynchronous protocol has achieved before. Designing RACS while achieving G1, G2 and G3 simultaneously is not trivial. First, designing a protocol that ensures robustness against adversarial network conditions without incurring additional overhead under synchronous conditions is a non-trivial algorithmic challenge. Second, designing a robust protocol while ensuring ease of understanding and compatibility with existing cloud software stacks is a challenge that no previous asynchronous consensus protocol has successfully addressed.
@@
I-A Contributions
@@
  • We provide the formal proofs of RACS in Appendix A.
@@
  • We show that RACS achieves nine times greater robustness than Raft under adversarial network conditions, while SADL-RACS delivers 150% higher throughput than Raft.
@@ II-A Threat Model and Assumptions
FLP [33] impossibility result claims that no deterministic protocol can guarantee both safety and liveness in a fully asynchronous system with even one faulty process. Practical protocols address the FLP [33] impossibility using two approaches: (1) assuming partial synchrony, and (2) using randomization. Protocols like Raft [14] and Multi-Paxos [13] assume partial synchrony and ensure safety at all times, however, lose liveness under asynchrony. Randomized protocols, in contrast, employ randomness instead of determinism, preserving safety always and ensuring liveness under asynchrony with probability 1.
In a typical cloud network, there are periods in which the network behaves synchronously, followed by phases where the network shows adversarial behavior. RACS makes use of this network behavior and operates in two modes: (1) synchronous mode and (2) fallback mode. During the synchronous periods, RACS employs a leader-based design to reach consensus using one round trip network delay, and during the fallback mode, RACS employs randomization.
@@ II-B Consensus
Multi-leader protocols [21, 20, 23, 24] and other extensions of leader-based protocols [35][36][22][37] also lose liveness under adversarial network conditions due to their reliance on view change mechanisms.
Randomized consensus protocols [27][26] overcome the adversarial network challenge by using randomization instead of view changes. However, they are rarely used in cloud applications due to (1) high performance overhead under synchronous network conditions, (2) incompatibility with the existing cloud software stack, and (3) complexity in understanding and implementing. In contrast, our protocol RACS provides a practical solution to adversarial network challenges by achieving: (1) robustness under adversarial network conditions, (2) optimal one round-trip commit latency under synchronous networks, and (3) a simple design for fast implementation and seamless industry adoption.
Rabia [38] and RACS take opposite paths on the simplicity–robustness tradeoff. Rabia [38] aims for simplicity in synchronous LANs but loses liveness under WAN asynchrony. In contrast, RACS ensures liveness under asynchrony. Rabia’s [38] leaderless design differs from Paxos/Raft standard leader-based model, requiring ground-up changes in the implementation, making adoption costly. Rabia [38] has quadratic message complexity, whereas RACS has linear message complexity, under synchrony.
QuePaxa [25] solves the asynchronous consensus challenge, however, significantly differs from RACS in design. QuePaxa [25] uses randomized local-coins and splits the replica role into separate proposer role and recorder role. In contrast, RACS employs global-coins and closely follows the standard Raft model.
Mahi-Mahi [39], Tusk [28], and Dag-Rider [40] are “blockchain” asynchronous protocols targeting Byzantine faults, which is out of our scope. We omit detailed comparisons due to space and relevance.
@@ III-C RACS Algorithm
Synchronous mode: The synchronous mode of RACS is leader-based and is similar to Raft, as depicted in lines 1–11 of Figure 4.. The synchronous mode leader $L_{vcur}$ for each view $v$ is predetermined and known to all replicas on bootstrap, for example $L_{vcur}=v_{cur}\%n$. As we discuss below, $L_{vcur}$ begins its tenure only after receiving a quorum of $<$vote$>$ messages, which allows $L_{vcur}$ to learn the most up-to-date block—a condition indirectly achieved in Raft [14] by electing the most up-to-date replica during view change.
@@
Fallback mode: The number of blocks proposed during the fallback mode is a tunable parameter that we explain in Section IV-A. For simplicity of illustration, we use 2 fallback blocks per view in the following description. Upon receiving $n/2+1$ $<$timeout$>$ messages, RACS enters the fallback mode of operation. In the fallback mode, all replicas act as leaders. Each replica takes the highest $block_{high}$ they are aware of, forms a level $1$ fallback block $B$ with a monotonically increasing rank compared to the highest $block_{high}$ it received and sends a $<$propose-async$>$ message (see line 12–16 of Figure 4).
@@
Once $n/2+1$ replicas have submitted $<$asynchronous-complete$>$ messages, the fallback mode is complete. The remaining task is to consistently select one chain from the set of fallback chains. This step is crucial for maintaining agreement. If two replicas select different fallback chains, the system loses the guarantee that all replicas commit the same sequence of blocks. To ensure that each replica consistently chooses the same chain, we employ a common-coin, an existing technique [26, 38] used in distributed algorithms.
@@
Upon receiving $n/2+1$ $<$asynchronous-complete$>$ messages, each replica individually flips a common-coin (see line 21 of Figure 4) to get the elected leader Lelected. Each replica commits a level $2$ fallback block $B$ from Lelected if $B$ arrived amongst the first $n/2+1$ $<$asynchronous-complete$>$ messages (see line 22 of Figure 4). If a replica observes that the level $2$ fallback block from Lelected does not appear in the first $n/2+1$ $<$asynchronous-complete$>$ messages, but appears in Bfall[Lelected], then, the replica sets blockhigh to Bfall[Lelected] (see line 23 of Figure 4). All replicas then exit the fallback mode and resume the synchronous path by uni-casting a $<$vote$>$ message to the synchronous leader Lvcur of the next view (see line 27 of Figure 4).
RACS fallback mode ensures liveness under adversarial network conditions; it only requires receiving at least $n/2+1$ messages in each asynchronous step. This quorum condition is trivially satisfied under the standard assumption that fewer than $n/2$ nodes can fail and with reliable point-to-point links (TCP). Thus, delayed messages or timeout settings do not impact the asynchronous liveness guarantee of RACS.
RACS protocol summary: Based on the above RACS explanation, its outcomes are summarized as follows. RACS ensures robustness against adversarial network conditions by committing new commands during the randomized fallback mode. RACS commits client commands in one network round trip, and hence provides optimal synchronous normal case latency. Finally, because RACS follows a simple design and extends Raft, it enables seamless integration with the existing cloud software stack. Although the RACS fallback mode modifies Raft [14] in non-trivial ways, the fallback mode employs the same replica and network structures, without requiring any additional cloud software stack changes. Thus, RACS remains fully compatible with existing cloud software stack and requires only incremental integration effort.
@@ III-D Correctness and Complexity
The formal proofs have been deferred Appendix A.
@@
Termination property: RACS ensures termination under both synchronous and adversarial network conditions, due to common-coin based randomization.
@@ IV-A Tuning fallback blocks
In Section III, we described RACS with two fallback blocks per view. In practice, fixed blocks per view are suboptimal under unpredictable adversarial conditions.
In RACS, we dynamically adjust the number of blocks in fallback mode using Multi-Armed Bandit techniques [44]. We bootstrap RACS with 2 blocks per fallback mode, and each replica monitors s: the number of times the algorithm switches between synchronous and fallback mode. A high s implies a network where adversarial conditions persist, whereas a low s indicates a transient network problem. RACS employs a Multi-Armed Bandit-based explore/exploit strategy, increasing the number of blocks per fallback mode when s increases and decreasing it when s decreases. RACS uses the same technique as Raft’s reconfiguration to consistently configure s across all replicas. Further details are omitted due to space.
@@ IV-B Application specific optimizations
For class 1 applications, we employ pipelining, a well-established technique in distributed algorithms [25, 13]. In pipelined-RACS, the synchronous mode leader $L_{\text{vcur}}$ proposes multiple blocks before receiving $n/2+1$ $<$vote$>$ messages, allowing multiple blocks to be committed within a single round-trip network latency.
IV-B1 Optimization for class 2 applications
@@
RACS blocks contain two types of data: (1) metadata about the block such as $v$, and $r$, and (2) a batch of client commands. Typically, the metadata is only a few bytes in size, while the batch of client commands can span several kilobytes. Bundling both metadata and client commands within RACS blocks increases the block size, which limits the system’s throughput due to the bandwidth bottleneck at the leader replica. This is a significant limitation of RACS, particularly for class 2 applications that demand high throughput.
To address this limitation, we propose SADL, a novel command dissemination layer that decouples command dissemination from the critical path of the consensus protocol. SADL is a consensus-agnostic command dissemination layer designed to reliably distribute client commands to at least a majority of replicas, ensuring efficient and fault-tolerant command propagation. SADL provides fault tolerance and availability for individual client batches, but does not guarantee a total order of commands. Once SADL replicates commands across a majority of replicas, RACS layer no longer needs to include the commands in it’s blocks. Instead, RACS only embeds a small, fixed-size ”reference” to the batch already disseminated by SADL. Using a fixed-size reference instead of the entire command batch significantly reduces the RACS leader replica’s bandwidth consumption, leading to a substantial improvement in throughput, as demonstrated in Section VI.
SADL design significantly differs from existing dissemination approaches. Unlike existing dissemination mechanisms [28, 48], where the dissemination layer advances in a lock-step fashion, SADL allows each replica to build its own chain of SADL blocks independently of the pace of other replicas. This key difference enables SADL to achieve higher throughput than existing approaches. We present this subtle yet non-trivial distinction as a novel systems contribution of this paper.
@@ V Implementation
We implemented pipelined-RACS and SADL-RACS using Go version 1.19 [29], in 3661 and 4631 lines of codes, respectively. We used the standard Go network library and Protobuf encoding [49]. We used a single-threaded event based design for RACS’s main event handling logic. All events—message receipt, timeouts, etc.—are handled by a single thread, ensuring deterministic event execution. For better I/O performance, we use multiple threads for TCP message reception, as done in prior work [38, 25, 43]. Both RACS and SADL-RACS implement batching in both clients and replicas. SADL-RACS implementation is publicly available on github [^1] https://github.com/ISTA-SPiDerS/sadl-racs.